# Adversarial Generation and Collaborative Evolution of Safety-Critical Scenarios for Autonomous Vehicles

### v1

@@ 1 Introduction
As autonomous vehicles (AVs) [17, 41, 37, 20] approach widespread deployment, ensuring their safety and reliability to earn public trust has become critical [25, 24, 21, 34]. However, this endeavor is challenging due to the scarcity of real-world data on rare yet critical incidents. Simulation-based testing [2, 36] offers a controlled, reusable, and cost-effective way of evaluating behavior under various conditions, particularly safety-critical scenarios that probe their safety capacity. Current scenario generation approaches, which largely rely on predefined threat templates [58, 56, 50] or rule-based strategies [6, 49, 5, 59], struggle to reveal the full spectrum of safety flaws. Due to these weak risk exposure abilities, AVs retain undiscovered vulnerabilities from inadequate validation. This challenge is closely related to the field of adversarial machine learning. In this domain, research focuses on exposing the vulnerabilities of deep neural networks through various attack methodologies. Such work also involves evaluating model robustness and developing more resilient architectures [48, 29, 33, 57, 44, 26, 23, 30, 28, 27, 35, 32, 15, 31, 22].
To address these limitations, we propose ScenGE, a two-stage framework that exposes safety vulnerabilities in AV by performing adversarial threat generation and collaborative trajectory evolution. Our first stage, Meta-Scenario Generation, uses a LLM to creatively generate a core adversarial threat from a benign text prompt. To ensure this threat is both plausible and challenging, a retrieval augmented generation RAG [53] framework grounds the LLM’s reasoning in a structured knowledge base of traffic regulations, driver qualification standards, and realistic pre-crash scenarios. The generated meta-scenario is then expressed as executable Scenic code [11, 12], allowing for diverse and scalable instantiation in the CARLA simulator [8]. However, the threat created by a single adversarial agent is often predictable and insufficient to create a truly critical dilemma for the AV. Our second stage, Complex Scenario Evolution, therefore crafts more complex threats by coordinating the surrounding background traffic. It first builds an adversarial collaborator graph to identify the most influential vehicles. The trajectories of these key agents are then carefully optimized. Rather than causing direct collisions or simple chaos, these optimizations intensify interaction complexity by strategically limiting the ego’s escape paths and obstructing its line-of-sight, ultimately increasing the likelihood of a critical incident.
@@ 2 Related Work
Simulation-Based Testing for AV. Simulation-based testing, facilitated by platforms such as CARLA [8], MetaDrive [18], and LimSim [52], offers a cost-effective and controlled environment for evaluating AVs under diverse driving conditions. A key advantage is the ability to replicate rare yet critical scenarios that are impractical or hazardous to test in the real world. This capability for controlled, repeatable testing is critical for systematically identifying performance vulnerabilities and ultimately validating the safety of AVs.
Safety-Critical Scenario Generation. The generation of safety-critical scenarios is crucial for evaluating AV safety, with existing approaches including generative models that learn from real-world data [6, 43, 42, 39, 9], optimization-based methods that synthesize scenarios via tailored objectives [4, 49, 3, 54], and semantic-driven methods that incorporate high-level context [14, 62, 50, 51, 61]. More recent works leverage language models for generation from text or logs [19, 60], or enhance coverage through semantic replay and trajectory compression [47, 10].
@@ 3.2 Meta-Scenario Generation
Safety Driving Knowledge Construction. The knowledge base $\mathbb{D}=\{\mathbf{D}_{r},\mathbf{D}_{l},\mathbf{D}_{c}\}$ consists of three components, each representing a distinct aspect of driving knowledge essential for simulating normative and adversarial traffic behavior. ❶ $\mathbf{D}_{r}$ contains 27 driving regulations segmented from official manuals in the USA, Germany, and China, covering behaviors such as lane merging, overtaking, and other maneuvers. ❷ $\mathbf{D}_{l}$ includes 100 standardized driver’s license test questions and answers that assess traffic rule knowledge, situational awareness, and safe behavior selection. Together, $\mathbf{D}_{r}$ and $\mathbf{D}_{l}$ provide normative behavioral priors intentionally violated to construct safety-critical adversarial behaviors. In contrast, ❸ $\mathbf{D}_{c}$ comprises 14 pre-crash scenarios drawn from taxonomies in the NHTSA Pre-Crash Typology Report [38] (e.g., unprotected left turns, red-light violations), providing concrete adversarial patterns for scenario construction. Collectively, these components inform the synthesis of plausible threat scenarios and support the generation of critical adversarial conditions.
LLM-Driven Scenario Generation. Given a simple prompt $\Phi_{\mathrm{base}}$ of a benign scene, the LLM is prompted to generate a detailed, safety-critical scenario by introducing one main adversarial agent $\mathbf{a}_{\mathrm{adv}}$ into the scenario. It infers the agent’s properties and the associated environmental context through in-context learning [7]. However, simply adopting an LLM may lead to unsafe or unrealistic critical scenarios; thus, we ground the reasoning process in structured driving knowledge. To this end, relevant knowledge is retrieved from the database $\mathbb{D}$ and combined with the instruction prompt $\Phi_{\mathrm{inst}}$ to form the input to the LLM $f_{\mathrm{LLM}}$. The generation process is formalized as:
$\begin{split}&\langle{\Phi}_{\mathrm{c}},{\Phi}_{\mathrm{p}},{\Phi}_{\mathrm{b}},{\Phi}_{\mathrm{R}},{\Phi}_{\mathrm{L}}\rangle=\\ &\qquad f_{\mathrm{LLM}}\left(\mathbf{a}_{\mathrm{ego}},\Phi_{\mathrm{base}},f_{R}(\mathbb{D},\Phi_{\mathrm{base}})\mid\Phi_{\mathrm{inst}}\right),\end{split}$ (3.1)
@@ 3.3 Complex Scenario Evolution
$\begin{split}\mathcal{L}&=\lambda_{1}\underbrace{\|\tilde{\tau}_{i^{\star}}-\tilde{\tau}_{\text{ego}}\|_{2}}_{\mathcal{L}_{\text{ego}}}+\lambda_{2}\underbrace{\|\Delta^{2}\tilde{\tau}_{i^{\star}}\|_{2}^{2}}_{\mathcal{L}_{\text{smooth}}}\\ &+\lambda_{3}\underbrace{\|(\tilde{\tau}_{i^{\star}}-\tilde{\tau}_{\text{ego}})\times(\tilde{\tau}_{\text{adv}}-\tilde{\tau}_{\text{ego}})\|_{\perp}}_{\mathcal{L}_{\text{occ}}}.\end{split}$ (3.4)
@@ 4 Experiment and Evaluation
4.1 Experimental Setup
Simulation environment and benchmark. We utilise the CARLA simulator [8], an open-source and highly customizable urban driving simulator, to create a closed-loop simulation environment. We adopt SafeBench [55] as the benchmarking framework, which supports diverse RL-based AV agents and standardized evaluation. Following [58], we use 8 base traffic scenarios (e.g., Straight Obstacle, Lane Changing) curated from the NHTSA Pre-Crash Typology Report [38], each containing 10 diverse driving routes. For each route, 10 adversarial scenarios are generated, resulting in 800 challenging scenarios for evaluation and comparison per method.
AV algorithms. Following [58], we mainly employ 3 representative RL-based AV algorithms as testing agents, including Proximal Policy Optimization (PPO) [40], Soft Actor-Critic (SAC) [16], and Twin Delayed Deep Deterministic Policy Gradient (TD3) [13].
Compared baselines. We compare our ScenGE with several existing scenario generation methods, including Learning-to-Collide (LC) [6], AdvSim (AS) [49], Carla Scenario Generator (CS) [5], Adversarial Trajectory Optimization (AT) [59], and ChatScene [58]. For fair comparisons, each method is applied on the same 8 base scenarios and routes to generate 800 challenging scenarios under consistent generation logic and evaluation settings.
Metrics. Following SafeBench [55], we adopt a set of key metrics to evaluate AV performance in generated scenarios. Two core indicators are used: the collision rate (CR $\uparrow$) measures the frequency of collisions and reflects safety risk, and the overall score (OS $\downarrow$) aggregates system-level performance. In addition, we evaluate three additional dimensions: the safety level (frequency of running red lights (RR $\uparrow$), frequency of running stop signs (SS $\uparrow$), and average distance driven out of road (OR $\uparrow$), the functionality level (route following stability (RF $\downarrow$), average percentage of route completion (Comp $\downarrow$), and average time spent to complete the route (TS $\uparrow$)), and the etiquette level (average acceleration (ACC $\uparrow$), average yaw velocity (YV $\uparrow$), and frequency of lane invasion (LI $\uparrow$). Higher ($\uparrow$) values indicate worse performance, while $\downarrow$ indicates the contrary.
Implementation details. All experiments were conducted on a server with an Intel(R) Core(TM) i9-14900K CPU and two NVIDIA GeForce RTX 4090 GPUs with 24GB memory. The LLM used for Meta-Scenario Generationis qwq-32b [46], the reasoning model from the Qwen series. In the Complex Scenario Evolution module, we construct 10 background vehicles and perturb the trajectories of 4 selected ones, each over 60% of their trajectory. The 4 perturbed vehicles are selected based on the highest attention relevance to ego and adversarial agents. The 60% perturbation window is centered around each vehicle’s most relevant keyframe. We set $\gamma$ in the decay matrix to $0.8$. In the loss calculation, we set $\lambda_{1}=0.3$, $\lambda_{2}=0.2$, and $\lambda_{3}=0.5$.
4.2 Main Results
@@
Straight
Obstacle
Turning
Obstacle
Lane
Changing
Vehicle
Passing
Red-light
Running
Unprotected
Left-turn
Right-
turn
Crossing
Negotiation
Avg.
@@ 4.3 Ablation Studies
Knowledge Prior. Removing $\mathbf{D}_{r}$ yields CR 79.4% and OS 52.3%, reflecting its role in guiding rule-focused violations. Removing $\mathbf{D}_{l}$ gives CR 77.4% and OS 55.2%, showing its effect on enhancing logical consistency in behavior. Removing $\mathbf{D}_{c}$ results in CR 80.5% and OS 52.1%, confirming its importance in producing realistic and high-risk scenarios.
LLM. Both GPT-4o and qwq-32b demonstrate strong capabilities, yielding similar CRs of 81.3% and 82.0%, respectively. This suggests that the performance is not critically sensitive to the specific choice of a powerful LLM.
Collaborator Graph. We ablate the number of perturbed agents by selecting 2, 4, and 6 collaborators. Perturbing 4 agents performs best with CR 82% and OS 49.9%, balancing adversarial strength and scenario plausibility. Interestingly, perturbing 6 agents (CR 78.1%) is less effective than perturbing 2 (CR 80.3%). We hypothesize that this occurs because an excessive number of agents transforms our intended coordinated, precise threat into easily avoidable chaos. Their mutual interference and abnormal behavior likely prompt the AV to adopt a conservative policy. This hypothesis is also supported by OS, where the 6-agent setting yields a higher score (56.1%) compared to the 2-agent setting (55.1%).
Temporal Modeling. The full setting (with both mask and decay) yields the best result with CR 82% and OS 49.9%. Removing the temporal mask reduces temporal causality in collaborator selection, leads to CR 79.3% and OS 55.6%, while removing the temporal decay results in CR 78.2% and OS 57.3%. These results highlight the complementary role of both components in capturing temporally coherent influence.
Perturbation Ratio. We compare three perturbation ratios centered around the selected keyframe: 50%, 60%, and 70%. Perturbing 60% of the segment achieves the best result with CR 82% and OS 49.9%. 50% leads to CR 80.2% and OS 53.4%, indicating insufficient behavioral deviation, while 70% causes CR 81.1% and OS 52.1% due to over-modification and reduced plausibility. These results shows moderate ratio balances realism and adversarial effect.
Loss. We ablate each component in $\mathcal{L}$ to assess its contribution. Removing $\mathcal{L}_{\text{ego}}$ leads to CR 78.6% and OS 55.3%, reflecting reduced collision targeting. Excluding $\mathcal{L}_{\text{smooth}}$ yields CR 80.2% and OS 54.3%, with trajectories becoming visibly unstable. Removing $\mathcal{L}_{\text{occ}}$ results in CR 79.4% and OS 56.8%, indicating weaker alignment between adversary and ego. The full loss yields the best trade-off, and ablating any term consistently reduces CR and increases OS.

### v2

@@ 1 Introduction
As autonomous vehicles (AVs) hu2023_uniad ; shao2024lmdrive ; ma2024dolphins ; li2022bevformer approach widespread deployment, ensuring their safety and reliability to earn public trust has become critical liang2022parallel ; liang2020efficient ; liang2022large ; liu2025natural . However, this endeavor is challenging due to the scarcity of real-world data on rare yet critical incidents. Simulation-based testing cai2025text2scenario ; lu2024multimodal offers a controlled, reusable, and cost-effective way of evaluating behavior under various conditions, particularly safety-critical scenarios that probe their safety capacity. Current scenario generation approaches, which largely rely on predefined threat templates zhang2024chatscene ; xu2025diffscene ; wang2024drivedreamer or rule-based strategies ding2020learning ; wang2021advsim ; scenariorunner ; zhang2022adversarial , struggle to reveal the full spectrum of safety flaws. Due to these weak risk exposure abilities, AVs retain undiscovered vulnerabilities from inadequate validation. This challenge is closely related to the field of adversarial machine learning. In this domain, research focuses on exposing the vulnerabilities of deep neural networks through various attack methodologies. Such work also involves evaluating model robustness and developing more resilient architectures wang2021dual ; liu2019perceptual ; liu2020bias ; zhang2021interpreting ; tang2021robustart ; liang2023badclip ; liang2024object ; liu2021training ; liu2020spatiotemporal ; liu2023x ; liu2022harnessing ; liu2023exploring ; guo2023towards ; liu2023towards ; liang2025revisiting .
@@
To address these limitations, we propose ScenGE, a two-stage framework that exposes safety vulnerabilities in AV by performing adversarial threat generation and collaborative trajectory evolution. Our first stage, Meta-Scenario Generation, uses a LLM to creatively generate a core adversarial threat from a benign text prompt. To ensure this threat is both plausible and challenging, a retrieval augmented generation RAG wu2024retrieval framework grounds the LLM’s reasoning in a structured knowledge base of traffic regulations, driver qualification standards, and realistic pre-crash scenarios. The generated meta-scenario is then expressed as executable Scenic code fremont2019scenic ; fremont2022scenic , allowing for diverse and scalable instantiation in the CARLA simulator Dosovitskiy17 . However, the threat created by a single adversarial agent is often predictable and insufficient to create a truly critical dilemma for the AV. Our second stage, Complex Scenario Evolution, therefore crafts more complex threats by coordinating the surrounding background traffic. It first builds an adversarial collaborator graph to identify the most influential vehicles. The trajectories of these key agents are then carefully optimized. Rather than causing direct collisions or simple chaos, these optimizations intensify interaction complexity by strategically limiting the ego’s escape paths and obstructing its line-of-sight, ultimately increasing the likelihood of a critical incident.
@@ 2 Related Work
Simulation-Based Testing for AV. Simulation-based testing, facilitated by platforms such as CARLA Dosovitskiy17 , MetaDrive li2021metadrive , and LimSim wen2023limsim , offers a cost-effective and controlled environment for evaluating AVs under diverse driving conditions. A key advantage is the ability to replicate rare yet critical scenarios that are impractical or hazardous to test in the real world. This capability for controlled, repeatable testing is critical for systematically identifying performance vulnerabilities and ultimately validating the safety of AVs.
Safety-Critical Scenario Generation. The generation of safety-critical scenarios is crucial for evaluating AV safety, with existing approaches including generative models that learn from real-world data ding2020learning ; tan2021scenegen ; suo2021trafficsim ; rempe2022generating ; feng2023trafficgen , optimization-based methods that synthesize scenarios via tailored objectives chen2021adversarial ; wang2021advsim ; cao2022advdo ; xiang2023v2xp , and semantic-driven methods that incorporate high-level context gao2023magicdrive ; zhong2023language ; wang2024drivedreamer ; wang2024driving ; Zheng2025OccWorld . More recent works leverage language models for generation from text or logs li2024chatsumo ; zhao2024chat2scenario , or enhance coverage through semantic replay and trajectory compression tian2024llm ; feng2023dense .
@@ 3.2 Meta-Scenario Generation
Safety Driving Knowledge Construction. The knowledge base $\mathbb{D}=\{\mathbf{D}_{r},\mathbf{D}_{l},\mathbf{D}_{c}\}$ consists of three components, each representing a distinct aspect of driving knowledge essential for simulating normative and adversarial traffic behavior. ❶ $\mathbf{D}_{r}$ contains 27 driving regulations segmented from official manuals in the USA, Germany, and China, covering behaviors such as lane merging, overtaking, and other maneuvers. ❷ $\mathbf{D}_{l}$ includes 100 standardized driver’s license test questions and answers that assess traffic rule knowledge, situational awareness, and safe behavior selection. Together, $\mathbf{D}_{r}$ and $\mathbf{D}_{l}$ provide normative behavioral priors intentionally violated to construct safety-critical adversarial behaviors. In contrast, ❸ $\mathbf{D}_{c}$ comprises 14 pre-crash scenarios drawn from taxonomies in the NHTSA Pre-Crash Typology Report najm2007pre (e.g., unprotected left turns, red-light violations), providing concrete adversarial patterns for scenario construction. Collectively, these components inform the synthesis of plausible threat scenarios and support the generation of critical adversarial conditions.
LLM-Driven Scenario Generation. Given a simple prompt $\Phi_{\mathrm{base}}$ of a benign scene, the LLM is prompted to generate a detailed, safety-critical scenario by introducing one main adversarial agent $\mathbf{a}_{\mathrm{adv}}$ into the scenario. It infers the agent’s properties and the associated environmental context through in-context learning dong2024survey . However, simply adopting an LLM may lead to unsafe or unrealistic critical scenarios; thus, we ground the reasoning process in structured driving knowledge. To this end, relevant knowledge is retrieved from the database $\mathbb{D}$ and combined with the instruction prompt $\Phi_{\mathrm{inst}}$ to form the input to the LLM $f_{\mathrm{LLM}}$. The generation process is formalized as:
$\langle{\Phi}_{\mathrm{c}},{\Phi}_{\mathrm{p}},{\Phi}_{\mathrm{b}},{\Phi}_{\mathrm{R}},{\Phi}_{\mathrm{L}}\rangle=f_{\mathrm{LLM}}\left(\mathbf{a}_{\mathrm{ego}},\Phi_{\mathrm{base}},f_{R}(\mathbb{D},\Phi_{\mathrm{base}})\mid\Phi_{\mathrm{inst}}\right),$ (3.1)
@@ 3.3 Complex Scenario Evolution
$\mathcal{L}$ $=\lambda_{1}\underbrace{\|\tilde{\tau}_{i^{\star}}-\tilde{\tau}_{\text{ego}}\|_{2}}_{\mathcal{L}_{\text{ego}}}+\lambda_{2}\underbrace{\|(\tilde{\tau}_{i^{\star}}-\tilde{\tau}_{\text{ego}})\times(\tilde{\tau}_{\text{adv}}-\tilde{\tau}_{\text{ego}})\|_{\perp}}_{\mathcal{L}_{\text{occ}}}+\lambda_{3}\underbrace{\|\Delta^{2}\tilde{\tau}_{i^{\star}}\|_{2}^{2}}_{\mathcal{L}_{\text{smooth}}}.$ (3.4)
@@ 4 Experiment and Evaluation
Straight Obstacle Turning Obstacle Lane Changing Vehicle Passing Red-light Running Unprotected Left-turn Right- turn Crossing Negotiation Avg.
@@
4.1 Experimental Setup
Simulation environment and benchmark. We utilise the CARLA simulator Dosovitskiy17 , an open-source and highly customizable urban driving simulator, to create a closed-loop simulation environment. We adopt SafeBench xu2022safebench as the benchmarking framework, which supports diverse RL-based AV agents and standardized evaluation. Following zhang2024chatscene , we use 8 base traffic scenarios (e.g., Straight Obstacle, Lane Changing) curated from the NHTSA Pre-Crash Typology Report najm2007pre , each containing 10 diverse driving routes. For each route, 10 adversarial scenarios are generated, resulting in 800 challenging scenarios for evaluation and comparison per method.
AV algorithms. Following zhang2024chatscene , we mainly employ 3 representative RL-based AV algorithms as testing agents, including Proximal Policy Optimization (PPO) schulman2017proximal , Soft Actor-Critic (SAC) haarnoja18b , and Twin Delayed Deep Deterministic Policy Gradient (TD3) fujimoto18a .
Compared baselines. We compare our ScenGE with several existing scenario generation methods, including Learning-to-Collide (LC) ding2020learning , AdvSim (AS) wang2021advsim , Carla Scenario Generator (CS) scenariorunner , Adversarial Trajectory Optimization (AT) zhang2022adversarial , and ChatScene zhang2024chatscene . For fair comparisons, each method is applied on the same 8 base scenarios and routes to generate 800 challenging scenarios under consistent generation logic and evaluation settings.
Metrics. Following SafeBench xu2022safebench , we adopt a set of key metrics to evaluate AV performance in generated scenarios. Two core indicators are used: the collision rate (CR $\uparrow$) measures the frequency of collisions and reflects safety risk, and the overall score (OS $\downarrow$) aggregates system-level performance. In addition, we evaluate three additional dimensions: the safety level (frequency of running red lights (RR $\uparrow$), frequency of running stop signs (SS $\uparrow$), and average distance driven out of road (OR $\uparrow$), the functionality level (route following stability (RF $\downarrow$), average percentage of route completion (Comp $\downarrow$), and average time spent to complete the route (TS $\uparrow$)), and the etiquette level (average acceleration (ACC $\uparrow$), average yaw velocity (YV $\uparrow$), and frequency of lane invasion (LI $\uparrow$). Higher ($\uparrow$) values indicate worse performance, while $\downarrow$ indicates the contrary.
Implementation details. All experiments were conducted on a server with an Intel(R) Core(TM) i9-14900K CPU and two NVIDIA GeForce RTX 4090 GPUs with 24GB memory. The LLM used for Meta-Scenario Generationis qwq-32b qwq32b , the reasoning model from the Qwen series. In the Complex Scenario Evolution module, we construct 10 background vehicles and perturb the trajectories of 4 selected ones, each over 60% of their trajectory. The 4 perturbed vehicles are selected based on the highest attention relevance to ego and adversarial agents. The 60% perturbation window is centered around each vehicle’s most relevant keyframe. We set $\gamma$ in the decay matrix to $0.8$. In the loss calculation, we set $\lambda_{1}=0.3$, $\lambda_{2}=0.2$, and $\lambda_{3}=0.5$.
4.2 Main Results
@@ 4.3 Ablation Studies
❶ Knowledge Prior. Removing $\mathbf{D}_{r}$ yields CR 79.4% and OS 52.3%, reflecting its role in guiding rule-focused violations. Removing $\mathbf{D}_{l}$ gives CR 77.4% and OS 55.2%, showing its effect on enhancing logical consistency in behavior. Removing $\mathbf{D}_{c}$ results in CR 80.5% and OS 52.1%, confirming its importance in producing realistic and high-risk scenarios.
❷ LLM. Both GPT-4o and qwq-32b demonstrate strong capabilities, yielding similar CRs of 81.3% and 82.0%, respectively. This suggests that the performance is not critically sensitive to the specific choice of a powerful LLM.
❸ Collaborator Graph. We ablate the number of perturbed agents by selecting 2, 4, and 6 collaborators. Perturbing 4 agents performs best with CR 82% and OS 49.9%, balancing adversarial strength and scenario plausibility. Interestingly, perturbing 6 agents (CR 78.1%) is less effective than perturbing 2 (CR 80.3%). We hypothesize that this occurs because an excessive number of agents transforms our intended coordinated, precise threat into easily avoidable chaos. Their mutual interference and abnormal behavior likely prompt the AV to adopt a conservative policy. This hypothesis is also supported by OS, where the 6-agent setting yields a higher score (56.1%) compared to the 2-agent setting (55.1%).
❹ Temporal Modeling. The full setting (with both mask and decay) yields the best result with CR 82% and OS 49.9%. Removing the temporal mask reduces temporal causality in collaborator selection, leads to CR 79.3% and OS 55.6%, while removing the temporal decay results in CR 78.2% and OS 57.3%. These results highlight the complementary role of both components in capturing temporally coherent influence.
❺ Perturbation Ratio. We compare three perturbation ratios centered around the selected keyframe: 50%, 60%, and 70%. Perturbing 60% of the segment achieves the best result with CR 82% and OS 49.9%. 50% leads to CR 80.2% and OS 53.4%, indicating insufficient behavioral deviation, while 70% causes CR 81.1% and OS 52.1% due to over-modification and reduced plausibility. These results shows moderate ratio balances realism and adversarial effect.
❻ Loss. We ablate each component in $\mathcal{L}$ to assess its contribution. Removing $\mathcal{L}_{\text{ego}}$ leads to CR 78.6% and OS 55.3%, reflecting reduced collision targeting. Excluding $\mathcal{L}_{\text{smooth}}$ yields CR 80.2% and OS 54.3%, with trajectories becoming visibly unstable. Removing $\mathcal{L}_{\text{occ}}$ results in CR 79.4% and OS 56.8%, indicating weaker alignment between adversary and ego. The full loss yields the best trade-off, and ablating any term consistently reduces CR and increases OS.